OpenStudy (anonymous):

s=n/2(a+t) solve for n

OpenStudy (anonymous):

the choices are?

OpenStudy (jadeishere):

s = n / 2a + 2t n = s * 2a + 2t

OpenStudy (anonymous):

thank ya